Wine is approximately 12% ethanol by volume. ethanol has a molar mass of 46.06 g/mol and a density of 0.789g/ml. How many moles of ethanol are present in a 750 ml of wine?


1
Expert's answer
2021-03-02T02:00:57-0500

The volume of ethanol in 750 ml of wine is: 750 x 0.12 = 90 ml

Density = mass/volume

Mass = density x volume

Mass of pure ethanol = 0.789 x 90 = 71.01 g

n = m/M

n (ethanol) = 71.01 / 46.06 = 1.54 mol